OpenStudy (anonymous):

integrate(from 0 to 1): (xe - xe^x)dx i dnt kno wat do with the e...

OpenStudy (helder_edwin):

e is a constant (unless there's a mistake) so you have \[ \large \int_0^1(xe-xe^x)\,dx=e\int_0^1x\,dx-\int_0^1xe^x\,dx \]

OpenStudy (anonymous):

yes y = e , its a line

OpenStudy (anonymous):

how do u do the second part?

OpenStudy (helder_edwin):

integration by parts \[\large \int udv=uv-\int vdu \] here \[ u=x\qquad dv=e^x\,dx \]

OpenStudy (anonymous):

ok thanks